# MMAU: A Holistic Benchmark of Agent Capabilities Across Diverse Domains

### v1

@@ Abstract
Recent advances in large language models (LLMs) have increased the demand for comprehensive benchmarks to evaluate their capabilities as human-like agents. Existing benchmarks, while useful, often focus on specific application scenarios, emphasizing task completion but failing to dissect the underlying skills that drive these outcomes. This lack of granularity makes it difficult to deeply discern where failures stem from. Additionally, setting up these environments requires considerable effort, and issues of unreliability and reproducibility sometimes arise, especially in interactive tasks. To address these limitations, we introduce the Massive Multitask Agent Understanding (MMAU) benchmark, featuring comprehensive offline tasks that eliminate the need for complex environment setups. It evaluates models across five domains, including Tool-use, Directed Acyclic Graph (DAG) QA, Data Science and Machine Learning coding, Contest-level programming and Mathematics, and covers five essential capabilities: Understanding, Reasoning, Planning, Problem-solving, and Self-correction. With a total of 20 meticulously designed tasks encompassing over 3K distinct prompts, MMAU provides a comprehensive framework for evaluating the strengths and limitations of LLM agents. By testing 18 representative models on MMAU, we provide deep and insightful analyses. Ultimately, MMAU not only sheds light on the capabilities and limitations of LLM agents but also enhances the interpretability of their performance. Datasets and evaluation scripts of MMAU are released at https://github.com/apple/axlearn/docs/research/mmau.
@@ 3.2.1 Tool-Use
We curated an in-house dataset for tool-use with conversation trajectories following the standard tool-use (a.k.a. function-calling) protocol. We select from a subset of RapidAPI Hub  [^1] https://rapidapi.com/hub functions and ask human annotators to create realistic scenarios with user queries, ground truth function calls and actual function returns from the RapidAPI endpoints. In total, our in-house tool-use dataset consists of 409 single-step (task:single-tool-use) and 258 multi-step tool-use conversations (task: multi-turn multi-tool-use). Out of the 409 single-step tool use conversations, 225 require making parallel tool calls (task: parallel-tool-use). Figure LABEL:fig:tool-use-examples shows some of these examples. We adapt this dataset for the following tasks.
@@
Task: DAG QA In this task, a user presents a set of requirements to which the LLM must respond by selecting and ordering a sequence of tool invocations from multiple choices provided. This design examines whether the model can identify the relevant tools and deduce the correct dependencies between them. The prompt enumerates the possible tool use orderings from which the LLM agent is asked to pick one, and the label is derived from the ground truth function call sequence. For example, we transform the multi-step example in Figure LABEL:fig:tool-use-examples into a DAG QA task as shown in Figure 14
@@ Acknowledgement
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix D
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Code, data, and instructions needed to reproduce the main experimental results are included in the supplemental material.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix D
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix A
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix D Data examples
Appendix E Datasheet for Datasets
We follow the Datasheet for Datasets framework to answer the important questions considering this dataset.
E.1 Motivation
For what purpose was the dataset created?    See Section 1.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?    Please refer to our author list for details.
Who funded the creation of the dataset?    The dataset incurred no monetary cost other than the time and effort of the authors.
E.2 Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?    See Section 3.2.
How many instances are there in total (of each type, if appropriate)?    See Table 4.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?    The full dataset is provided.
Is there a label or target associated with each instance?    See Section 3.2 for the definitions of each data type and the associated target.
Is any information missing from individual instances?    No.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?    No.
Are there recommended data splits?    The dataset is entirely intended for evaluation purpose.
Are there any errors, sources of noise, or redundancies in the dataset?    We have carefully reviewed each data sample to ensure the reliability of the benchmark. We are committed to actively improving and enriching the benchmark based on community feedback.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?    Yes, the dataset is self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?    No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?    No.
Does the dataset identify any subpopulations (e.g., by age, gender)?    No.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset?    No.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?    No.
E.3 Collection Process
How was the data associated with each instance acquired?    See Section 3.2.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?    See Section 3.2.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?    N/A.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?    The authors volunteered to perform the necessary manual data curation.
Over what timeframe was the data collected?    From April 2024 to Jun 2024.
Were any ethical review processes conducted (e.g., by an institutional review board)?    We carefully reviewed our data curation procedure. While we believe the source datasets permit our use herein, we do not make any representations as to the ethical review process of the original dataset. Please follow the terms of such datasets for further consideration.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?    See Section 3.2.
Were the individuals in question notified about the data collection?    N/A.
Did the individuals in question consent to the collection and use of their data?    N/A.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?    N/A.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?    N/A.
E.4 Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?    See Section 3.2.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?    See Section 3.2.
Is the software that was used to preprocess/clean/label the data available?    N/A.
E.5 Uses
Has the dataset been used for any tasks already?    See Section 3.2.
Is there a repository that links to any or all papers or systems that use the dataset?    We will host the benchmark data within the Hugging Face community, which has convenient tracking features on the reference and usage of the dataset.
What (other) tasks could the dataset be used for?    The dataset can be generally customized and used for evaluating the capabilities of large language models and large multimodal models.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?    No.
Are there tasks for which the dataset should not be used?    N/A.
E.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?    The dataset will be fully open-sourced.
How will the dataset will be distributed (e.g., tarball on the website, API, GitHub)?    The dataset will be made available as a Hugging Face dataset.
When will the dataset be distributed?    Upon publishing the manuscript.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?    No.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?    No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?    No.
E.7 Maintenance
Who will be supporting/hosting/maintaining the dataset?    The authors will be responsible for supporting/hosting/maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?    Please refer to our author list for details.
Is there an erratum?    No.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?    Yes.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., was the individuals in question were told that their data would be retained for a fixed period of time and then deleted)?    N/A.
Will older versions of the dataset continue to be supported/hosted/maintained?    Yes.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?    Yes. We welcome contributions via the Hugging Face community.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have increased the demand for comprehensive benchmarks to evaluate their capabilities as human-like agents. Existing benchmarks, while useful, often focus on specific application scenarios, emphasizing task completion but failing to dissect the underlying skills that drive these outcomes. This lack of granularity makes it difficult to deeply discern where failures stem from. Additionally, setting up these environments requires considerable effort, and issues of unreliability and reproducibility sometimes arise, especially in interactive tasks. To address these limitations, we introduce the Massive Multitask Agent Understanding (MMAU) benchmark, featuring comprehensive offline tasks that eliminate the need for complex environment setups. It evaluates models across five domains, including Tool-use, Directed Acyclic Graph (DAG) QA, Data Science and Machine Learning coding, Contest-level programming and Mathematics, and covers five essential capabilities: Understanding, Reasoning, Planning, Problem-solving, and Self-correction. With a total of 20 meticulously designed tasks encompassing over 3K distinct prompts, MMAU provides a comprehensive framework for evaluating the strengths and limitations of LLM agents. By testing 18 representative models on MMAU, we provide deep and insightful analyses. Ultimately, MMAU not only sheds light on the capabilities and limitations of LLM agents but also enhances the interpretability of their performance. Datasets and evaluation scripts of MMAU are released at https://github.com/apple/axlearn/blob/main/docs/research/mmau.
@@ 3.2.1 Tool-Use
We curated an in-house dataset for tool-use with conversation trajectories following the standard tool-use (a.k.a. function-calling) protocol. We select from a subset of RapidAPI Hub  [^1] https://rapidapi.com/hub functions and ask human annotators to create realistic scenarios with user queries, ground truth function calls and actual function returns from the RapidAPI endpoints. In total, our in-house tool-use dataset consists of 409 single-step (task:single-tool-use) and 258 multi-step tool-use conversations (task: multi-turn multi-tool-use). Out of the 409 single-step tool use conversations, 225 require making parallel tool calls (task: parallel-tool-use). Figs. 10, 11 and 12 show some of these examples. We adapt this dataset for the following tasks.
@@
Task: DAG QA In this task, a user presents a set of requirements to which the LLM must respond by selecting and ordering a sequence of tool invocations from multiple choices provided. This design examines whether the model can identify the relevant tools and deduce the correct dependencies between them. The prompt enumerates the possible tool use orderings from which the LLM agent is asked to pick one, and the label is derived from the ground truth function call sequence. For example, we transform the multi-step example in Figure 11 into a DAG QA task as shown in Figure 14